OpenStudy (anonymous):

Solve for x. |2x + 2| = 14 x = 6 and x = -6 x = 6 and x = -8 x = 8 and x = -8 x = -6 and x = 8

OpenStudy (jiteshmeghwal9):

|2x+2|=14 =>2x+2=14 =>2x=14-2 =>2x=12 =>x=12/2=6.

OpenStudy (anonymous):

im confused it is asking for two answers

OpenStudy (jiteshmeghwal9):

If there is positive digit in the the bracket & if there is negative the answer will come as negative . now, we don't that the digit in the bracket is negative or positive . so, the answer will come as both positive or negative.i.e 6 & -6.

OpenStudy (anonymous):

THANKS

jimthompson5910 (jim_thompson5910):

The other answer isn't -6. Here's why |2x + 2| = 14 |2(-6) + 2| = 14 |-12+2| = 14 |-10| = 14 10 = 14 ... which is false So x = -6 isn't a solution